1.Find the Domain and Range of the following functions if 𝑓: 𝑅 β†’ 𝑅 i) 𝑓(π‘₯) = 12 + π‘™π‘œπ‘” (1 βˆ’ π‘₯^2 ) ii) 𝑓(π‘₯) = π‘₯^2 +1/π‘₯^2 +5

2. Find if the functions are surjective, injective and bijective if 𝑓: 𝑅 β†’ 𝑅 i) 𝑓(π‘₯) = π‘₯^5 + 5 ii) 𝑓(π‘₯) = 𝑒^4π‘₯Β 


1
Expert's answer
2022-03-03T13:19:21-0500

1.

i)


"f(x)=12+\log(1-x^2)"

"1-x^2>0=>-1<x<1"

Domain: "(-1, 1)"


"0<1-x^2\le1"

Range: "(-\infin, 12]."


ii)


"f(x)=x^2+\dfrac{1}{x^2}+5"

"x\not=0"

Domain: "(-\infin, 0)\cup (0, \infin)"


"x^2+\dfrac{1}{x^2}\ge2, x\in\R"

Range: "[7, \infin)."


2)

i)


"f(x)=x^5+5"

"f(x)" is strictly increasing onΒ "\R." So, ifΒ "n, m\in \R, n\not=m," then "n>m" or "n<m."

Without loss of generality, assume thatΒ "n>m."

Then we haveΒ "n^5+5>m^5+5." So if "n\not=m," then "f(n)\not=f(m)."

Therefore the function "f(x)=x^5+5" is injective.


"f(x)" is strictly increasing onΒ "\R." For every "\forall y\in \R, \exist x\in \R" such that


"f(x)=y"

"x^5+5=y=>x=(y-5)^{1\/5}"

Therefore the function "f(x)=x^5+5" is surjective.


"f(x)" is strictly increasing onΒ "\R." For every "\forall y\in \R, \exist! x\in \R" such that


"f(x)=y"

"x^5+5=y=>x=(y-5)^{1\/5}"

Therefore the function "f(x)=x^5+5" is bijective.


ii)

"\ud835\udc53(\ud835\udc65) = \ud835\udc52^{4\ud835\udc65 }"

"f(x)" is strictly increasing onΒ "\R." So, ifΒ "n, m\in \R, n\not=m," then "n>m" or "n<m."

Without loss of generality, assume thatΒ "n>m."

Then we haveΒ "e^{4n}>e^{4m}" . So if "n\not=m," then "f(n)\not=f(m)."

Therefore the function "f(x)=e^{4x}" is injective.


If "y\le0," we cannot find "x\in \R" such that "f(x)=y."

Therefore the function "f(x)=e^{4x}" is not surjective and is not bijective.